OpenStudy (kainui):

Here's a fun problem I came across. During an hour two independent events can happen at any time. What's the probability that the events are at least 10 minutes apart?

OpenStudy (rational):



OpenStudy (kainui):

Hahaha yeah you got it.

OpenStudy (rational):

P(|E1-E2| > 10 minutes) = (Area of shaded region) / (Total area) = 5^2/6^2 = 25/36 ?

OpenStudy (kainui):

Yep, exactly. =)

OpenStudy (rational):

these are really fun xD i got some practice a couple of months ago when @amistre64 was preparing for some probability theory exam

OpenStudy (kainui):

Something that just occurred to me that I don't know the answer to is what is the probability of having 3 events 10 minutes apart? Or, is there a good reason why the probability is 5^2/6^2? If the events were separated by 20 minutes would the answer then become 4^2/6^2 ? I guess this would be easy to check for the general answer.

OpenStudy (rational):

very interesting, 3 events requires a triple integral is it

OpenStudy (rational):

second part of the question is easy yeah we always get squares in top and bottom because of symmetry : |E1-E2| > 20

OpenStudy (kainui):

for the 2D case in general the answer is: \[\frac{(60-t)^2}{60^2}\] which is quite nice!

Parth (parthkohli):

Wow, that's beautiful.

Parth (parthkohli):

Looks like my book has this... is this what is called "infinitistic probability"?

OpenStudy (kainui):

Yup, I guess? I guess continuous probability distribution maybe? It's important for quantum mechanics!

OpenStudy (ikram002p):

:)

OpenStudy (dan815):